# On the Design and Analysis of LLM-Based Algorithms

### v2

@@ 4.3.3 Insights
    LLM nodes responsible for reasoning and answering only occupy a small fraction of the costs of the overall algorithm, while playing a critical role in the output of the algorithm and thus in the error metrics. Therefore, our general recommendation is to prompt these LLM calls to think step by step before answering, instead of answering directly, which will most likely boost the accuracy significantly with negligible loss in efficiency.
@@
4.4 Experiments
We validate our analysis via experiments with a Qwen-2-72B-Instruct model [YYH+24], supported by vLLM [KLZ+23] and running on a server with 4 Nvidia A100-80G GPUs. For each experiment, we vary the depth, width or degree of the grid of variables (which controls the difficulty of task instances), while validating Insight 1 by comparing two options for prompting LLM calls that are responsible for reasoning and answering, or Insight 2 by comparing the “cyclic” and “parallel” options for retrieval.
Error metrics of interest include (1) the exact-match error, which takes value 00 if the answer returned by the algorithm matches the ground-truth solution exactly, and value $1$ otherwise; (2) the absolute error, i.e. the absolute value of the difference between the algorithm’s answer and the ground-truth solution; and (3) the missed-coverage error, i.e. the ratio of needles that the algorithm fails to retrieve and hence are not included in the references maintained by the algorithm. Note that missed coverage might arise not just from failures of the LLM calls responsible for retrieval, but also from the possibility that the algorithm terminates too early with fewer rounds than necessary, due to hallucination by the LLM calls responsible for reasoning and answering. Cost metrics of interest are the same as those explained in Section 3.2, plus the number of rounds, a metric specific to the hierarchical pattern of algorithms considered in this section.
@@ 4.4 Experiments
Empirical results for this experiment are illustrated in Figure 13, which confirm Insight 1, i.e. prompting the LLM to think step by step for reasoning and answering significantly boosts the accuracy of the overall algorithm while incurring minor overhead in cost metrics, compared to answering directly. Note from the last row, though, that the relative difference in terms of end-to-end latencies increases with the parallelism degree $p$, which has been predicted by our analysis, in particular Eq. (6).
@@
Empirical results for these two experiments can be found in Figures 14 and 15, which confirm that the considered algorithm achieves high accuracy for a wide range of task configurations. Insight 2 from our analysis is also validated: compared with the “cyclic” version, the “parallel” version incurs a larger number of rounds and hence higher cost metrics, except for the end-to-end latencies with parallelism, for which the benefits of parallelism outweight the downside of requiring more rounds.
@@
5 Case study: recursive decomposition
This section studies LLM-based algorithms with recursive task decomposition, a pattern that is vastly different from those studied in previous sections, and yet still covered by the analytical framework proposed in Section 2. Roughly speaking, a recursive LLM-based algorithm starts from the original task of concern and recursively generates more intermediate sub-tasks, so that each sub-task can be solved by aggregating the solutions to its children tasks, while being unaware of other sub-tasks involved in the overall algorithm. In particular, solving and/or decomposing each sub-task can be achieved by LLM calls, while the outline of recursive task decomposition remains symbolic. Such LLM-based algorithms have been widely applied, e.g. in LAMBADA [KKB+23], LLM programs [SSC+23], ADaPT [PKH+24], THREAD [SMLG24], Recursion of Thought [LK23], Decomposed Prompting [KTF+23], ReDel [ZDCB24], among many others.
In the following, we first introduce the concrete task under consideration, and then design a recursive LLM-based algorithm for solving it, highlighting the dynamic construction of its computational graph at runtime. We then provide analysis for its accuracy and efficiency using the proposed framework, which is further validated with numerical experiments; along the way, we derive a formal guarantee for generic LLM-based algorithms, under the technical assumption of additive errors and bounded sensitivity.
5.1 Concrete setup
For concreteness, we consider the same task introduced in Section 4.1 and visualized in Figure 11, which is about calculating a targeted variable based on clues about a grid of many variables.
One major difference here is that we consider much larger depth $d$, width $w$ and/or degree $g$ for the grid of variables, which implies a larger DAG of needles, i.e. clues relevant to the targeted question. Even if all relevant clues are given a priori, the complex reasoning required to answer the targeted question correctly can be well beyond the capability of one single LLM call, which motivates decomposing the reasoning process into multiple LLM calls.
The other major difference is, we assume that the clues are not directly given in plain text, but rather need to be accessed via querying a database. For each query, the database takes a name as input, and returns a clue for the variable of the same name, e.g. “A2 = B1 + B3” for input “A2” or “C1 = 10” for input “C1”. The LLM-based algorithm need to decide by itself what to query from the database. Such a setting is motivated by, and can be regarded as an abstraction of, practical scenarios where an autonomous agent in the wild need to actively retrieve relevant information by itself in order to accomplish a task (unlike in typical problem-solving benchmarks where such information is given), via querying a real database or knowledge graph, using a search engine, retrieving from documents, etc. In this case study, we assume that querying the database does not involve LLM calls, and thus neglect its costs in our analysis.
5.2 Algorithm
@@
    Otherwise, query the database with the input variable, and invoke one LLM call with the returned clue, which is prompted to either answer with a numeric value or conduct further task decomposition, i.e. identifying variables whose values will be necessary and sufficient for calculating the input variable.
@@
We make a few comments about this recursive algorithm. (1) The computational graph of this algorithm, or a generic LLM-based algorithm following the pattern of recursive decomposition, is constructed dynamically in a depth-first-search style at runtime. See Figure 16 for a visualization. The final graph is symmetric about the nodes corresponding to the leaf variables (D, E); on the left are LLM calls for identifying children variables for the non-leaf variables (A, B, C), while on the right are LLM calls for calculating their numeric values. (2) The process of complex reasoning is decomposed recursively, so that each LLM call within this algorithm only need to handle a small step of reasoning, whose complexity is irrelevant to that of the overall task. (3) One particular configuration of this algorithm, like in the previous case study, is about how the LLM calls (especially the ones responsible for calculation) are prompted, e.g. to answer directly or think step by step. The impact of this design choice on the accuracy and efficiency of the overall algorithm will be investigated analytically and empirically soon in this section.
Figure 16: The computational graph of a recursive LLM-based algorithm is constructed dynamically, while a dictionary of solved tasks is maintained (to avoid solving the same intermediate sub-task repetitively) throughout one run of the algorithm.
5.3 Analysis
5.3.1 Error metrics
@@
Recall from Eq. (1) that for a specific node $v$ with inputs $x_{1},\dots,x_{k}$, the error of its output $y$ is assumed to be upper bounded by $\mathcal{E}(y)\leq f_{v}(\mathcal{E}(x_{1}),\mathcal{E}(x_{2}),\dots,\mathcal{E}(x_{k}))$ for some function $f_{v}$. For the concrete setting of the current case study, there are two major failure modes for the aforementioned algorithm: errors in recursive task decomposition, and errors in calculating the numeric values of specific variables. Empirical results during our early exploration suggested that LLMs make very few mistakes of the first kind, while the second failure mode is much more common, due to the limited arithmetic capabilities of current LLMs. Moreover, all mathematical operations considered in this concrete settings (including addition, substraction, maximum and minimum) are $1$-Lipschitz continuous with respect to each input variable.
All these motivate us to derive the following error bound, which indeed holds true for any generic LLM-based algorithm represented by a DAG, under the assumption of additive errors and bounded sensitivity. The proof is deferred to Section 5.5 after the experiments.
Proposition 2.
Suppose that the assumption of additive errors and bounded sensitivity holds true for an LLM-based algorithm represented by a DAG, namely for each node $v$ with $k$ inputs $x_{1},\dots,x_{k}$ and a single output $y$, it holds that
for some node-specific additive error $\mathcal{E}_{v}$ and finite sensitivity parameter $S\geq 0$. Then the error of the output $y(v)$ of any node $v$, including the one that generates the final output of the overall algorithm, is bounded by
$\mathcal{E}(y(v))\leq\sum_{w\in\text{DAG}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow v)}S^{|\text{path}|}\times\mathcal{E}_{w}.$ (7)
[…]
This proposition precisely characterizes how the error $\mathcal{E}_{w}$ of each node $w$ impacts the accuracy of the overall algorithm, via the number and lengths of the paths from each node to the final output.
5.3.2 Cost metrics
@@
Let us first consider the cost of each LLM call. For each leaf variable, one LLM call is needed for finding its numeric value based on its corresponding clue of length $O(1)$, which implies $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(1)$ and $\mathcal{L}_{\textsf{dec}}\leq O(1)$, and thus $\mathcal{C}(\text{leaf})\leq\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+1,1)$. For each non-leaf variable, two LLM calls are needed. One of them is responsible for task decomposition, i.e. identifying its $g$ children variables, which implies $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(g)$ and $\mathcal{L}_{\textsf{dec}}\leq O(g)$, and thus $\mathcal{C}(\text{decomposition})\leq\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+g,g)$. The other LLM call is responsible for calculating the numeric value of the current variable based on the values of its children variables, which has $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(g)$, and $\mathcal{L}_{\textsf{dec}}\leq O(1)$ if the LLM is prompted to answer directly, or $\mathcal{L}_{\textsf{dec}}\leq\mathcal{L}_{\textsf{calc}}(g)$ for some function $\mathcal{L}_{\textsf{calc}}$ if it is prompted to do the calculation step by step (e.g. $\mathcal{L}_{\textsf{calc}}(g)\leq O(g)$ or $O(g^{2})$, depending on how the LLM executes the calculation with $g$ variables). In sum, we have $\mathcal{C}(\text{calculation})\leq\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+g,1\text{ or }\mathcal{L}_{\textsf{calc}}(g))$ for each LLM call responsible for calculation.
@@
The analysis above has also revealed the total number of LLM calls. Let $n$ and $n_{\textsf{leaf}}$ be the number of relevant variables and the number of relevant leaf variables, respectively. These parameters are determined by $d,w,g$ and other factors, as is the case for the $\ell$ parameter (total length of relevant clues) in Section 4. Recall from above that each leaf variable requires one LLM call for finding its value from its corresponding clue, while each non-leaf variable requires two LLM calls, one for task decomposition and one for calculation. Assuming that task decomposition is done correctly by the LLM for each non-leaf variable, the total number of LLM calls will be $2\times(n-n_{\textsf{leaf}})+n_{\textsf{leaf}}=2\times n-n_{\textsf{leaf}}$.
@@
In particular, prompting the LLM to do the calculation step by step, which is anticipated to significantly improve the accuracy of the overall algorithm, will increase most cost metrics (though not for the number of LLM calls or the total number of prefilling tokens) by a multiplicative factor, in contrast to the previous case study in Section 4 where step-by-step prompting incurs a minor additive cost.
5.4 Experiments
We validate our analysis via experiments with a Qwen-2-72B-Instruct model [YYH+24], supported by vLLM [KLZ+23] and running on a server with 4 Nvidia A100-80G GPUs. For each experiment, we vary the difficulty of the task through the depth $d$ or degree $g$ of the grid of variables, while comparing two options of prompting the LLM calls responsible for calculation, namely “answer directly” and “calculate step by step”.
@@ 5.4 Experiments
Empirical results are shown in Figures 17 and 18. For the former, we fix the width $w=6$ and degree $g=4$, while varying the depth $d$; for the latter, we fix the depth $d=3$ and width $w=100$, while varying the degree $g$. The results confirm our previous analysis: compared to “answer directly”, prompting the LLM calls to “calculate step by step” significantly boosts the accuracy of the overall algorithm and leads to satisfactory performance for a wide range of task configurations, while incurring higher costs in the latency and total number of decoding tokens by a multiplicative factor, and making no difference in the number of LLM calls or total number of prefilling tokens.
@@
Figure 17: Empirical results (10 trials) for reasoning with recursive task decomposition, where two options of prompting the LLM calls responsible for calculation are compared. The width $w=6$ and degree $g=4$ are fixed, while the depth $d$ varies.
@@
Figure 18: Empirical results (10 trials) for reasoning with recursive task decomposition, where two options of prompting the LLM calls responsible for calculation are compared. The depth $d=3$ and width $w=100$ are fixed, while the degree $g$ varies.
5.5 Proof of Proposition 2
$\mathcal{E}(y(v))\leq\mathcal{E}_{v}=\sum_{w\in\text{DAG}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow v)}S^{|\text{path}|}\times\mathcal{E}_{w}.$
$\mathcal{E}(y(u_{i}))\leq\sum_{w\in\text{DAG}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|}\times\mathcal{E}_{w},\quad i\in[k].$
$\leq\mathcal{E}_{v}+S\times\sum_{i\in[k]}\sum_{w\in\text{DAG}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|}\times\mathcal{E}_{w}$
$=\mathcal{E}_{v}+S\times\sum_{i\in[k]}\sum_{w\in\text{DAG}\setminus\{v\}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|}\times\mathcal{E}_{w}$
$=\mathcal{E}_{v}+S\times\sum_{w\in\text{DAG}\setminus\{v\}}\sum_{i\in[k]}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|}\times\mathcal{E}_{w}$
$=\mathcal{E}_{v}+\sum_{w\in\text{DAG}\setminus\{v\}}\sum_{i\in[k]}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|+1}\times\mathcal{E}_{w}.$
Here in the third line, we replace the summation over $w\in\text{DAG}$ with $w\in\text{DAG}\setminus\{v\}$, since there is certainly no path from node $v$ to its predecessor node $u_{i}$. To move forward, notice that $\mathcal{P}(w\rightarrow v)$ is a union of $k$ disjoint sets, where the $i$-th set contains paths ending with the directed edge from $u_{i}$ to $v$:
With this, we can further simplify the previous upper bound:
$\mathcal{E}(y(v))$ $\leq\mathcal{E}_{v}+\sum_{w\in\text{DAG}\setminus\{v\}}\sum_{i\in[k]}\sum_{\text{path}\in\mathcal{P}(w\rightarrow u_{i})}S^{|\text{path}|+1}\times\mathcal{E}_{w}$
$=\mathcal{E}_{v}+\sum_{w\in\text{DAG}\setminus\{v\}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow v)}S^{|\text{path}|}\times\mathcal{E}_{w}$
$=\sum_{w\in\text{DAG}}\sum_{\text{path}\in\mathcal{P}(w\rightarrow v)}S^{|\text{path}|}\times\mathcal{E}_{w}.$
This concludes our proof of the proposition.
6 Related works
LLM-based algorithms.
The concept of LLM-based algorithms, as was explained in Section 1, is fairly broad and general, ranging from a combination of one or multiple LLM calls with some prompt engineering [YYZ+23, BBK+23, BMZ+24, LLLD23, SLC+23, ZSH+23, WWS+23, PKH+24, CLGI24], to LLM-powered agent systems [MDL+23, WBZ+23, HZC+24, GLP+24, SST+23, ZWK+24, PKB+24, XCG+23, QXW+24, KSS+24] and compound AI systems [ZKC+24, CDH+24] that augment LLMs with additional abilities like tool use and long-term memory, and to the emerging paradigm of LLM programming [SSC+23, KTF+23, KSM+24, ZYX+24, KVG+24]. Some elements of our analytical framework proposed in Section 2, such as the principle of task decomposition, the computational-graph representation, evaluation of and comparison between LLM-based algorithms with accuracy and efficiency taken into account simultaneously, etc., have already appeared in one way or another in these prior works. It is primarily our unified, systematic and formal investigation into the design and analysis of generic LLM-based algorithms that distinguishes the current work from this vast literature.
Scaling properties of LLM test-time computation.
Recent works have started to investigate, analytically or empirically, the scaling properties of LLM test-time computation.
One line of research is concerned about the scaling properties of repeated sampling, e.g. randomly sampling multiple generations for the same prompt and then aggregating the results [CDH+24, BJE+24, SLXK24, WSL+24], which might be regarded as a stochastic version of parallel decomposition investigated in Section 3. Our work is orthogonal and complementary to this line of research, as our analysis is deterministic and targeted at generic patterns of LLM-based algorithms. Indeed, one potential direction for expanding the analytical framework proposed in this work would be to augment it with stochastic decoding and repeated sampling, along with relevant theoretical results from prior works.
Another line of works is concerned about improving the output quality of an LLM call, albeit at a higher cost, by generating more tokens autoregressively, e.g. via chains of thoughts [WWS+22] or step-by-step reasoning [KGR+22]. This idea has been investigated theoretically [FZG+23, MS24, LLZM24], and further popularized recently by the OpenAI o1 model [Ope24c]. As this approach of scaling up the test-time computation of individual LLM calls is becoming more widely adopted, it is important to understand, analytically and quantitatively, how it will impact the overall accuracy and efficiency when such LLM calls are embedded within an LLM-based algorithm. Our work can be useful in achieving this, as illustrated in the case studies in Sections 4 and 5, where the impacts of prompting LLM calls to “think step by step” versus “answer directly” can be characterized analytically with our proposed framework.
Learning-augmented algorithms.
Our work draws inspiration from the research area of algorithms with predictions / learning-augmented algorithms [MV22, LM22]. One standard paradigm of this area is to consider a specific task (say sorting [BC23] or clustering [EFS+22]), assume access to a black-box machine learning model that satisfies certain properties (e.g. what additional computation or information it can offer), propose a novel algorithm that leverages this ML model, provide theoretical guarantees for its accuracy and efficiency, and show the improvements over traditional, purely symbolic algorithms. Our work is similar in spirit to this line of research, but also substantially different, in that our analytical framework is targeted at general tasks and LLM-based algorithms, with assumptions on the capabilities, limitations and inference service of LLMs (regarded as general-purpose problem solvers) that are quite different from typical assumptions for task-specific ML models in the literature of learning-augmented algorithms.
7 Conclusion and discussion
We have introduced a formal framework for studying the design and analysis of LLM-based algorithms. After identifying the computational-graph representation, the principle of task decomposition, and other abstractions, we find it feasible to provide formal analysis for the accuracy and efficiency of generic LLM-based algorithms. In multiple case studies, we provide in-depth analysis with numerical validation for various patterns of LLM-based algorithms, including parallel, hierarchical and recursive decomposition.
Moving forward, we find it promising future research to further expand or apply the proposed framework and thereby advance the field of LLM-based algorithms, from a theoretical or practical perspective. For example:
  • We have been assuming that LLM nodes are state-less, and only considering the simplest and most straightforward usage of LLMs, i.e. sending a prompt and receiving a response for each LLM call. One potential extension is to consider LLM nodes with states, like short-term memory in agent systems, or other advanced ways of using LLMs, and understand how accuracy and efficiency of LLM-based algorithms are impacted in such cases.
  • In our empirical study, we have mainly focused on the hyperparameter $m$ that indicates the granularity of parallel task decomposition, or the option of prompting certain LLM calls to either answer directly or think step by step. Future work might place more emphasis on other configurations of LLM-based algorithms, such as the choices of LLM models (e.g. choosing a strong-but-expensive model for specific LLM calls and a weak-but-cheap one for the others within the same algorithm), LLM decoding methods (e.g. greedy vs. stochastic [WWS+23, CDH+24, BJE+24]), etc., and better understand how they impact the accuracy and efficiency of LLM-based algorithms.
  • Through our abstractions of multiple error and cost metrics for the same task, we have touched upon the topics of multi-objective optimization and hyperparameter optimization. Future work might try to formally investigate these aspects.
  • From a more practical perspective, future work might adopt the proposed framework, or some components and methodology within it, to assist the design, analysis, improvement and application of new LLM-based algorithms, or for fair comparison between different algorithms, with both accuracy and efficiency taken into account.
Given the plenty of opportunities for future research on LLM-based algorithms, we would like to invite the community to contribute to this exciting and rapidly developing field, using the current work as a starting point.
Acknowledgements
The authors would like to thank Yuchang Sun for helpful discussion, as well as the AgentScope team, especially Xuchen Pan, Dawei Gao and Weirui Kuang, for their assistance with using AgentScope and setting up LLM inference service for the experiments.

### v3

@@ C.3.3 Insights
    LLM nodes responsible for reasoning and answering only occupy a small fraction of the costs of the overall algorithm, while playing a critical role in the output of the algorithm and thus in the error metrics. Therefore, our general recommendation is to prompt these LLM calls to think step by step before answering (instead of answering directly), which will most likely boost the accuracy significantly with negligible loss in efficiency.
@@
C.4 Experiments
We validate our analysis via experiments with Qwen-2-72B-Instruct (Yang et al., 2024), supported by vLLM (Kwon et al., 2023) and running on a server with 4 Nvidia A100-80G GPUs. For each experiment, we vary the depth, width or degree of the grid of variables (which controls the difficulty of task instances), while validating the derived insights by comparing two options for prompting LLM calls that are responsible for reasoning and answering, or comparing the “cyclic” and “parallel” options for retrieval.
Error metrics of interest include (1) the exact-match error, which takes value 0 if the answer returned by the algorithm matches the ground-truth solution exactly, and value $1$ otherwise; (2) the absolute error, i.e., the absolute value of the difference between the algorithm’s answer and the ground-truth solution; and (3) the missed-coverage error, i.e., the ratio of needles that the algorithm fails to retrieve and hence are not included in the references maintained by the algorithm. Note that missed coverage might arise not just from failures of the LLM calls responsible for retrieval, but also from the possibility that the algorithm terminates too early with fewer rounds than necessary, due to hallucination by the LLM calls responsible for reasoning and answering. Cost metrics of interest are the same as those investigated in previous examples, plus the number of rounds, a metric specific to this case study.
@@ C.4 Experiments
Empirical results for this experiment are illustrated in Figure 17, which confirm our first insight, i.e., prompting the LLM to think step by step for reasoning and answering significantly boosts the accuracy of the overall algorithm while incurring minor overhead in cost metrics, compared to answering directly. Note from the last row, though, that the relative difference in terms of end-to-end latencies increases with the parallelism degree $p$, which has been predicted by our analysis, in particular Eq. (10).
@@
Empirical results for these two experiments can be found in Figures 18 and 19, which confirm that the considered algorithm achieves high accuracy for a wide range of task configurations. The second insight from our analysis is also validated: compared with the “cyclic” version, the “parallel” version incurs a larger number of rounds and hence higher cost metrics, except for the end-to-end latencies with parallelism, for which the benefits of parallelism outweigh the downside of requiring more rounds.
@@
Appendix D Supplementary materials for Section 4.3
Recursive LLM-based algorithms have been widely applied in prior works, e.g., in LAMBADA (Kazemi et al., 2023), LLM programs (Schlag et al., 2023), ADaPT (Prasad et al., 2024), THREAD (Schroeder et al., 2024), Recursion of Thought (Lee & Kim, 2023), Decomposed Prompting (Khot et al., 2023), ReDel (Zhu et al., 2024), among many others. This appendix includes full details for our case study, presented in Section 4.3, of recursive task decomposition. In the following, we first introduce the concrete task under consideration, and then design a recursive LLM-based algorithm for solving it, highlighting the dynamic construction of its computational graph at runtime. We then provide analysis for its accuracy and efficiency using the proposed framework, which is further validated with numerical experiments.
D.1 Concrete setup
For concreteness, we consider the same task introduced in Section C.1 and visualized in Figure 16, which is about calculating a targeted variable based on clues about a grid of many variables.
One major difference here is that we consider much larger depth $d$, width $w$ and/or degree $g$ for the grid of variables, which implies a larger DAG of needles, i.e., clues relevant to the targeted question. Even if all relevant clues are given a priori, the complex reasoning required to answer the targeted question correctly can be well beyond the capability of one single LLM call, which motivates decomposing the reasoning process into multiple LLM calls.
The other major difference is, we assume that the clues are not directly given in plain text, but rather need to be accessed via querying a database (or more generally, a source of information). For each query, the database takes a name as input, and returns a clue for the variable of the same name, e.g., “A2 = B1 + B3” for input “A2” or “C1 = 10” for input “C1”. The LLM-based algorithm need to decide by itself what to query from the database. Such a setting is motivated by, and can be regarded as an abstraction of, practical scenarios where an autonomous agent in the wild need to actively retrieve relevant information by itself in order to accomplish a task (unlike in typical problem-solving benchmarks where such information is given), via querying a real database or knowledge graph, using a search engine, retrieving from documents, etc. In this case study, we assume that querying the database does not involve LLM calls, and neglect its costs in our analysis for simplicity.
D.2 Algorithm
@@
    Otherwise, query the database with the input variable, and invoke one LLM call with the returned clue, which is prompted to either answer with a numeric value or conduct further task decomposition, i.e., identifying variables whose values will be necessary and sufficient for calculating the input variable.
@@
We make a few comments about this recursive algorithm. (1) The computational graph of this algorithm, or a generic LLM-based algorithm following the pattern of recursive decomposition, is constructed dynamically in a depth-first-search style at runtime. See Figure 7 for a visualization. The final graph is symmetric about the nodes corresponding to the leaf variables (D, E); on the left are LLM calls for identifying children variables for the non-leaf variables (A, B, C), while on the right are LLM calls for calculating their numeric values. (2) The process of complex reasoning is decomposed recursively, so that each LLM call within this algorithm only need to handle a small step of reasoning, whose complexity is irrelevant to that of the overall task. (3) One particular configuration of this algorithm, like in the previous case study, is about how the LLM calls (especially the ones responsible for calculation) are prompted, e.g., to answer directly or think step by step. The impact of this design choice on the accuracy and efficiency of the overall algorithm will be investigated analytically and empirically soon in this section.
D.3 Error analysis
@@
Recall from Eq. (1) that for a specific node $v$ with inputs $x_{1},\dots,x_{k}$, the error of its output $y$ is assumed to be upper bounded by $\mathcal{E}(y)\leq f_{v}(\mathcal{E}(x_{1}),\mathcal{E}(x_{2}),\dots,\mathcal{E}(x_{k}))$ for some function $f_{v}$. For the concrete setting of the current case study, there are two major failure modes for the aforementioned algorithm: errors in recursive task decomposition, and errors in calculating the numeric values of specific variables. Empirical results during our early exploration suggested that LLMs make very few mistakes of the first kind, while the second failure mode is much more common, due to the limited arithmetic capabilities of the LLMs under consideration. Moreover, all mathematical operations considered in this concrete settings (including addition, substraction, maximum and minimum) are $1$-Lipschitz continuous with respect to each input variable. Hence Proposition 2, which provides error bounds under the assumption of additive errors and bounded sensitivity, can be adopted for this case study, although we omit the details here for brevity.
D.4 Cost analysis
@@
Let us first consider the cost of each LLM call. For each leaf variable, one LLM call is needed for finding its numeric value based on its corresponding clue of length $O(1)$, which implies $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(1)$ and $\mathcal{L}_{\textsf{dec}}\leq O(1)$, and thus $\mathcal{C}(\text{leaf})\leq\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+1,1)$. For each non-leaf variable, two LLM calls are needed. One of them is responsible for task decomposition, i.e., identifying its $g$ children variables, which implies $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(g)$ and $\mathcal{L}_{\textsf{dec}}\leq O(g)$, and thus $\mathcal{C}(\text{decomposition})\leq\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+g,g)$. The other LLM call is responsible for calculating the numeric value of the current variable based on the values of its children variables, which has $\mathcal{L}_{\textsf{pre}}\leq\mathcal{L}_{\textsf{sys}}+O(g)$, and $\mathcal{L}_{\textsf{dec}}\leq O(1)$ if the LLM is prompted to answer directly, or $\mathcal{L}_{\textsf{dec}}\leq\mathcal{L}_{\textsf{calc}}(g)$ for some function $\mathcal{L}_{\textsf{calc}}$ if it is prompted to do the calculation step by step (e.g., $\mathcal{L}_{\textsf{calc}}(g)\leq O(g)$ or $O(g^{2})$, depending on how the LLM executes the calculation with $g$ variables). In sum, we have $\mathcal{C}(\text{calculation})\leq\mathcal{C}_{\textsf{LLM}}(\mathcal{L}_{\textsf{sys}}+g,1\text{ or }\mathcal{L}_{\textsf{calc}}(g))$ for each LLM call responsible for calculation.
@@
The analysis above has also revealed the total number of LLM calls. Let $n$ and $n_{\textsf{leaf}}$ be the number of relevant variables and the number of relevant leaf variables, respectively. These parameters are determined by $d,w,g$ and other factors, as is the case for the $\ell$ parameter (total length of relevant clues) in Section C. Recall from above that each leaf variable requires one LLM call for finding its value from its corresponding clue, while each non-leaf variable requires two LLM calls, one for task decomposition and one for calculation. Assuming that task decomposition is done correctly by the LLM for each non-leaf variable, the total number of LLM calls will be $2\times(n-n_{\textsf{leaf}})+n_{\textsf{leaf}}=2\times n-n_{\textsf{leaf}}$.
@@
In particular, prompting the LLM to do the calculation step by step, which is anticipated to significantly improve the accuracy of the overall algorithm, will increase most cost metrics (though not for the number of LLM calls or the total number of prefilling tokens) by a multiplicative factor, in contrast to the previous case study in Section C where step-by-step prompting incurs a minor additive cost overhead.
D.5 Experiments
@@
We validate our analysis via experiments with Qwen-2-72B-Instruct (Yang et al., 2024). For each experiment, we vary the task complexity while comparing two options of prompting the LLM calls responsible for calculation, namely “answer directly” and “calculate step by step”. Empirical results are shown in Figures 20 and 21. For the former, we fix the width $w=6$ and degree $g=4$, while varying the depth $d$; for the latter, we fix $d=3$ and $w=100$, while varying $g$. The results confirm our previous analysis: compared to “answer directly”, prompting the LLM calls to “calculate step by step” significantly boosts the accuracy of the overall algorithm and leads to satisfactory performance for a wide range of task configurations, while incurring higher costs in the latency and total number of decoding tokens by a multiplicative factor, and making no difference in the number of LLM calls or total number of prefilling tokens.
@@
Figure 20: Empirical results (10 trials) for reasoning with recursive task decomposition, where two options of prompting are compared. The width $w=6$ and degree $g=4$ are fixed, while the depth $d$ varies.
@@
Figure 21: Empirical results (10 trials) for reasoning with recursive task decomposition, where two options of prompting are compared. The depth $d=3$ and width $w=100$ are fixed, while the degree $g$ varies.